# Neural Network-Based Bandit: A Medium Access Control for the IIoT Alarm Scenario

### v1

@@ I-C ORGANIZATION
The paper is structured as follows. Section  II introduces the system model and problem formulation. Section  III describes the proposed NNBB scheme and the Medium Access Control (MAC) procedure followed by an IIoT device. We analyze the performance of NNBB through simulations in Section  IV. Finally, we conclude the paper in Section  V and also highlight some open research areas.
Notation: $\operatorname{argmax}(\cdot)$ and $\max(\cdot)$ are the argument of the maximum function and the maximum function, respectively. $\mathcal{P}(\cdot)$ is the power set and ||$\cdot$|| ⋅ | represents the cardinality of a set. Superscript $[\cdot]^{T}$ is the transpose operation and ${\|\cdot\|}_{2}$ is the Euclidean norm of a vector. Column vectors/matrices are denoted by boldface lowercase/uppercase letters, and $\mathbf{I}_{M}$ is the $M\times M$ identity matrix. $\mathbb{C}^{M\times 1}$ is a set of complex vectors of dimension $M\times 1$. A circularly-symmetric complex Gaussian random vector with mean $\mathbf{\bar{y}}$ and covariance matrix $\mathbf{Z}$ is represented by $\mathbf{y}\sim\mathcal{CN}(\mathbf{\bar{y}},\mathbf{Z})$. Furthermore, $\mathcal{U}(0,1)$ is a uniform distribution between 00 and $1$. Table II lists the symbols used in this paper and their definitions.
@@ II SYSTEM MODEL AND PROBLEM FORMULATION
Consider the IIoT alarm scenario in Fig. 1, where $M$ orthogonal channels are shared by a set $\mathcal{N}$ of wireless IIoT devices to communicate with the BS and the ExC. By utilizing $M$ orthogonal channels, the devices can transmit information concurrently over multiple channels. This significantly improves the likelihood of successful reception of the information at the receiver, in contrast to the scenario in which the devices have access to only a single channel. We assume that ${|\mathcal{N}|\gg M}$ and the devices are time synchronized [6]. At a certain time, an alarm is generated at a random location, called the epicenter. The devices that detect the alarm are called active devices, while the other devices are called inactive devices. Triggered by an alarm event, a random batch of devices ${\mathcal{N}^{\prime}\subseteq\mathcal{N}}$ becomes active. Let ${f(d_{\upsilon})}$ denote the activation probability function of device ${\upsilon\in\mathcal{N}}$. It is a decreasing function of the distance $d_{\upsilon}$ between the device $\upsilon$ and the alarm epicenter [4, 27].
The sole purpose of the active devices is to convey the alarm information to the ExC, which is not necessarily collocated with the BS, as illustrated in Fig. 1. Thus, they must successfully transmit the alarm information on at least one channel, irrespective of which active device makes it possible. For transmitting the alarm information, the active device ${\upsilon\in\mathcal{N}^{\prime}}$ chooses a transmission pattern denoted by ${\boldsymbol{a}_{\upsilon}=[a_{\upsilon,1},a_{\upsilon,2},\cdots,a_{\upsilon,M}]^{T}}$, where
@@
Figure 2: Illustration of the proposed protocol exemplified for the case of $M\!=$ 2 orthogonal channels. First, each active device transmits a pilot signal to the BS. Then, the BS broadcasts the received pilot signal to the active devices (context broadcast). Finally, the active devices transmit the alarm message after choosing their respective transmission patterns. In the structure of the alarm message, the alarm flag indicates the occurrence of an event, the header carries the metadata, and the payload contains relevant information for the ExC.
With $M$ channels, an active device can choose from $2^{M}$ transmission patterns. We compile all possible transmission patterns into a matrix ${\mathbf{\mathring{A}}=[\boldsymbol{\mathring{a}}_{1},\boldsymbol{\mathring{a}}_{2},\cdots,\boldsymbol{\mathring{a}}_{2^{M}}]\in\{0,1\}^{M\times 2^{M}}}$, where each column represents a unique transmission pattern and $\boldsymbol{\mathring{a}}_{i}\neq\boldsymbol{\mathring{a}}_{j},\forall i\neq j$. For instance, consider the case shown in Fig. 2 where $M=2$, $|\mathcal{N}^{\prime}|=5$, devices $1$, $2$, $4$ select channel $2$, device $5$ selects both channels, and device $3$ is remains silent. Then,
@@
Let us define the probability of the active device $\upsilon$ choosing a transmission pattern $\boldsymbol{\mathring{a}}_{i}$ as $\psi_{\upsilon}(\boldsymbol{\mathring{a}}_{i})$. Let the matrix ${\mathbf{\Psi}\in[0,1]^{|\mathcal{N}|\times 2^{M}}}$ store the elements of the following set ${\{\psi_{\upsilon}(\boldsymbol{\mathring{a}}_{i})|\forall i=1,\cdots,2^{M}}\}$ in its $\upsilon^{th}$ row. Additionally, ${\sum_{i=1}^{2^{M}}\psi_{\upsilon}(\boldsymbol{\mathring{a}}_{i})=1\textrm{,}\ \forall\upsilon\in\mathcal{N}^{\prime}}$. Now, the expected probability of a successful transmission can be written as [6]
$\!\!\!\Lambda(\mathbf{\Psi})=\!\!\!\!\!\!\sum_{\mathcal{N}^{\prime}\in\mathcal{P}(\mathcal{N})}\prod_{\upsilon\in\mathcal{N}^{\prime}}\!\!f(d_{\upsilon})\!\!\!\!\!\!\!\!\!\!\!\!\sum_{\mathbf{A}\in\{0,1\}^{M\times|\mathcal{N}^{\prime}|}}\!\!\!\!\!\!\!\!\!\!\!\Bigl{(}\xi(\mathcal{N}^{\prime},\mathbf{A})\!\prod_{\upsilon\in\mathcal{N}^{\prime}}\!\!\psi_{\upsilon}(\boldsymbol{a}_{\upsilon})\!\Bigl{)}.$ (8)
@@
$\begin{split}\underset{\mathbf{\Psi}}{\textrm{maximize}}\ &\Lambda\\ \textrm{subject to}\ &\sum_{i=1}^{2^{M}}\psi_{\upsilon}(\boldsymbol{\mathring{a}}_{i})=1\textrm{,}\ \forall\upsilon\in\mathcal{N}^{\prime},\\ &\mathbf{\Psi}\in[0,1]^{|\mathcal{N}|\times 2^{M}}.\end{split}$ (9)
@@ III NEURAL NETWORK-BASED BANDIT
NNBB is a DRL-based RA algorithm that is implemented in each IIoT device. The NNBB components are: (i) an agent (device) that receives a context, (ii) actions (transmission patterns), (iii) an action value ($\psi$) corresponding to each action, and (iv) a reward (feedback signal from the ExC) for updating the action values. The proposed protocol for the alarm transmission is illustrated in Fig. 2. The aforementioned NNBB components and the algorithm implementation are explained in detail below.
@@ III-A CONTEXT
An active agent (device) $\upsilon$ observes the context $\boldsymbol{s}_{\upsilon}$. For this, each active agent transmits a pilot signal consisting of $M$ pilot symbols, one on each channel. The BS then receives the signal aggregating all the pilot signals from the active devices and broadcasts it to the agents in the next time slot. The signal $\boldsymbol{s}_{\upsilon}$ received by the active agent $\upsilon$ informs it that some devices are active in the surrounding environment and implicitly captures hidden information about their identity. The goal of the NNBB-based RA mechanism is to make distributed access decisions based on such a context and learning experience.
@@
where $\rho$ represents the average transmit signal-to-noise ratio, ${\boldsymbol{\varphi}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})\in\mathbb{C}^{M\times 1}}$ is the Additive White Gaussian Noise (AWGN) with normalized power, and ${\mathbf{c}_{\upsilon}\in\mathbb{C}^{M\times 1}}$ captures the $M$ channel coefficients between the device $\upsilon$ and the BS. After receiving $\boldsymbol{s}$, the BS broadcasts $\boldsymbol{s}$ to the devices. Therefore, the context $\boldsymbol{s}_{\upsilon}$ received by the active device $\upsilon$ can be written as
@@
where ${\boldsymbol{\hat{\varphi}}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})\in\mathbb{C}^{M\times 1}}$ is the AWGN with normalized power. Here, we assume that $\rho$ is the same as the one used for pilot transmission to simplify the analysis without any loss of generality.
@@ III-B ACTION
After receiving $\boldsymbol{s}_{\upsilon}$, the active agent $\upsilon$ selects a transmission pattern $\boldsymbol{a}_{\upsilon}$, which we shall refer to as its action, and performs a single-hop alarm transmission to the ExC. Recall that with $M$ channels, an active device can choose from $2^{M}$ actions, which are collected as columns of the matrix $\boldsymbol{\mathring{A}}$. The alarm message consists of a header, alarm flag, and payload as shown in Fig. 2. The alarm flag indicates the occurrence of an event, the header carries the metadata, and the payload contains relevant information about the alarm for the ExC.
@@ III-D ACTION VALUE AND SELECTION
An agent aims to maximize its total reward obtained over time. To achieve this, it is attractive to choose those actions that have yielded a reward in the past. This is known as exploitation. However, to discover such reward-accruing actions, an agent must try new actions, which is known as exploration. In our work, an agent balances exploration and exploitation using the $\varepsilon$-greedy method. Such a method relies on the action value ${q(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i}),\ \forall i=1,\cdots,2^{M}}$ that represents a prediction of the expected reward for agent $\upsilon$ when it takes an action $\boldsymbol{\mathring{a}}_{i}$ with context $\boldsymbol{s}_{\upsilon}$. In the $\varepsilon$-greedy method, an agent $\upsilon$ primarily selects a greedy action (exploitation) ${\boldsymbol{a}_{\upsilon}=\operatorname{argmax}_{\boldsymbol{\mathring{a}}_{i}\in\{0,1\}^{M}}q(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i})}$, with occasional exploration by randomly selecting an action $\boldsymbol{a}_{\upsilon}$ from all possible actions with a probability $\varepsilon$. In this way, every action is eventually selected many times. It is necessary to decrease $\varepsilon$ to gradually emphasize exploitation over exploration. We gradually decrease $\varepsilon$ from $1$ to $0.1$ with a step size of $0.005$ after every alarm event.
@@
Due to channel fading and noise, the number of possible contexts is infinite, leading to an agent encountering a new context every time it becomes active. Thus, maintaining separate action values for each context is not a reasonable approach. Rather, the agent should maintain the action values as a parameterized function and tune its parameters to better match $q$ with the reward obtained after observing the feedback from the ExC. The parameterized function, for which a DNN is employed, allows an estimation of the action value without maintaining a table of action values.
@@
The parameterized action-value function for the agent $\upsilon$ is expressed as ${\hat{q}(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i},\mathbf{w})}$, where $\mathbf{w}$ is a vector of connection weights present in the DNN layers. The input to the DNN is the received context, as illustrated in Fig. 3. The DNN comprises two fully connected hidden layers and a dense output layer. Notice that two hidden layers are capable of approximating any smooth mapping with arbitrary accuracy [28]. Also, to grapple with the exploding gradients, we use gradient norm clipping [29]. The outputs from the DNN are $2^{M}$ parameterized action values $\hat{q}$. In addition, the DNN employs (a) ReLU non-linearity, which is a computationally efficient thresholding operation, compared to sigmoid and tanh, that helps alleviate the problem of vanishing gradients, and (b) RMSProp optimizer that maintains a moving average of the squared gradient values for each weight to adjust the update size for each weight. Moreover, the learning rate of the DNN is a hyper-parameter in RMSProp; we decrease it with a decay rate of $0.015$ per alarm event.
@@
After calculating the reward ${r(\boldsymbol{a}_{\upsilon})}$, the active agent $\upsilon$ stores the tuple ${\{\boldsymbol{s}_{\upsilon},\boldsymbol{a}_{\upsilon},r(\boldsymbol{a}_{\upsilon})\}}$ in its memory buffer $E$. If $E$ is full, it removes an already stored tuple from $E$ in a First-In-First-Out (FIFO) fashion. Indeed, there is no need to store the entire dataset in the agent’s memory.
@@
using RMSProp. Herein, $r_{j}$ and $\hat{q}_{j}$ represent the reward and parameterized action value, respectively, for the $j^{th}$ sample in the mini-batch. Meanwhile, the gradient vector $\nabla_{\mathbf{w}}\Omega$ is clipped in the following manner:
@@
III-E ALGORITHM IMPLEMENTATION
  1. 1.
    Transmit their respective pilot signals to the BS, which then broadcasts the received aggregated signal back to them. Consequently, each active agent acquires its context.
  2. 2.
    Feed their context to their DNN, which yields $2^{M}$ action values.
  3. 3.
  4. 4.
    Monitor the feedback from the ExC and determine its reward.
  5. 5.
  6. 6.
    Train their DNN.
Algorithm 1 provides a comprehensive description of NNBB.

### v2

@@ I-C ORGANIZATION
The paper is structured as follows. Section II introduces the system model and problem formulation. Section III describes the proposed NNBB scheme and the Medium Access Control (MAC) procedure followed by an IIoT device. Section IV discusses the computational complexity and convergence of NNBB, while Section V discusses the benchmark schemes. We analyze the performance of NNBB through simulations in Section VI. Finally, we conclude the paper in Section VII and also highlight some open research areas.
Notation: $\operatorname{argmax}(\cdot)$, $\max(\cdot)$, and $\min(\cdot)$ are the argument of the maximum function, the maximum function, and the minimum function, respectively. $\mathcal{P}(\cdot)$ is the power set and ||$\cdot$|| ⋅ | represents the cardinality of a set. Superscripts $[\cdot]^{T}$ and $[\cdot]^{H}$ are transpose and conjugate transpose operations, respectively, while ${\|\cdot\|}_{2}$ is the Euclidean norm of a vector. Column vectors/matrices are denoted by boldface lowercase/uppercase letters. $\textrm{Re}(\boldsymbol{x})$ and $\textrm{Im}(\boldsymbol{x})$ return column vectors consisting of the real and imaginary components, respectively, of $\boldsymbol{x}$, while $\textrm{abs}(\boldsymbol{x})$ returns a column vector with the absolute value of every element of $\boldsymbol{x}$. The ${M\times M}$ identity matrix and the ${M\times 1}$ null vector are denoted as $\mathbf{I}_{M}$ and ${\mathbf{0}_{M}}$, respectively. ${\mathbb{C}^{M\times 1}}$ and ${\mathbb{R}^{M\times 1}}$ are sets of complex and real vectors of dimension $M\times 1$, respectively, while ${\operatorname{diag}(\mathbf{X})}$ is a column vector comprising the diagonal elements of $\mathbf{X}$. A circularly-symmetric complex Gaussian random vector with mean $\mathbf{\bar{y}}$ and covariance matrix $\mathbf{Z}$ is represented by ${\mathbf{y}\sim\mathcal{CN}(\mathbf{\bar{y}},\mathbf{Z})}$. A uniform distribution between 00 and $1$ is denoted as ${\mathcal{U}(0,1)}$. Furthermore, ${\textrm{Pr}(X=x)}$ denotes the probability that the random variable $X$ takes on the value $x$. Table II lists the symbols used in this paper and their definitions.
@@ II SYSTEM MODEL AND PROBLEM FORMULATION
Consider the IIoT alarm scenario in Fig. 1, wherein a set $\mathcal{N}$ of wireless IIoT devices communicate with the BS and the ExC. Specifically, IIoT devices transmit data periodically from the sensed industrial process to the BS, which consolidates it before sending it onward for processing or storage. They are also continuously monitoring/sensing for alarm events, which they must inform to the the ExC. In this work, we focus on the timely alarm transmission process.
There are $M$ orthogonal channels for the IIoT devices to transmit alarm information concurrently. This significantly improves the likelihood of successful reception of the information at the receiver, in contrast to the scenario in which devices have access to only a single channel. We assume that ${|\mathcal{N}|\gg M}$ and devices are time synchronized [6]. At a certain time, an alarm is generated at a random location, called the epicenter. Several examples of an alarm event are available in Section I.I-B. Devices that detect the alarm are called active devices, while the other devices are called inactive devices. Triggered by an alarm event, a random batch of devices ${\mathcal{N}^{\prime}\subseteq\mathcal{N}}$ becomes active. Let ${f(d_{\upsilon})}$ denote the activation probability function of device ${\upsilon\in\mathcal{N}}$. It is a decreasing function of the distance $d_{\upsilon}$ between the device $\upsilon$ and the alarm epicenter [4, 32].
The sole purpose of the active devices is to convey the alarm information to the ExC, which is not necessarily collocated with the BS, as illustrated in Fig. 1. The ExC responds to the received alarm information by initiating control actions to manage the industrial process. Thus, active devices must successfully transmit the alarm information on at least one channel, irrespective of which active device makes it possible.
For transmitting the alarm information, the active device ${\upsilon\in\mathcal{N}^{\prime}}$ chooses a transmission pattern denoted by ${\boldsymbol{a}_{\upsilon}=[a_{\upsilon,1},a_{\upsilon,2},\cdots,a_{\upsilon,M}]^{T}}$, where
@@
Figure 2: Illustration of the proposed protocol exemplified for the case of $M\!=$ 2 orthogonal channels. First, each active device transmits a pilot signal to the BS. Then, the BS broadcasts the received pilot signal to devices (context broadcast). Finally, active devices transmit the alarm message after choosing their respective transmission patterns. In the structure of the alarm message, the alarm flag indicates the occurrence of an event, the header carries the metadata, and the payload contains relevant information for the ExC.
With $M$ channels, an active device can choose from $2^{M}$ transmission patterns. We compile all possible transmission patterns into a matrix ${\mathbf{\mathring{A}}=[\boldsymbol{\mathring{a}}_{1},\boldsymbol{\mathring{a}}_{2},\cdots,\boldsymbol{\mathring{a}}_{2^{M}}]\in\{0,1\}^{M\times 2^{M}}}$, where each column represents a unique transmission pattern and $\boldsymbol{\mathring{a}}_{i}\neq\boldsymbol{\mathring{a}}_{j},\forall i\neq j$. For instance, consider the case shown in Fig. 2 where $M=2$ and $|\mathcal{N}^{\prime}|=5$. Devices $1$, $2$, and $4$ select channel $2$, device $5$ selects both channels, and device $3$ remains silent. Then,
@@
Let us define the probability of the active device $\upsilon$ choosing a transmission pattern $\boldsymbol{\mathring{a}}_{i}$ as $\psi_{\upsilon}(\boldsymbol{\mathring{a}}_{i})$. Let the matrix ${\mathbf{\Psi}\in[0,1]^{|\mathcal{N}|\times 2^{M}}}$ store the elements of the set ${\{\psi_{\upsilon}(\boldsymbol{\mathring{a}}_{i})|\forall i=1,\cdots,2^{M}\}}$ in its $\upsilon^{th}$ row. Additionally, ${\sum_{i=1}^{2^{M}}\psi_{\upsilon}(\boldsymbol{\mathring{a}}_{i})=1\textrm{,}\ \forall\upsilon\in\mathcal{N}^{\prime}}$. Now, the probability of a successful transmission can be written as [6]
$\!\!\!\!\Lambda(\mathbf{\Psi})=\!\!\!\!\!\!\!\sum_{\mathcal{N}^{\prime}\in\mathcal{P}(\mathcal{N})}\overbrace{\prod_{\upsilon\in\mathcal{N}^{\prime}}\!\!f(d_{\upsilon})}^{g_{1}(\mathcal{N}^{\prime})}\!\!\!\!\!\!\!\!\!\!\!\!\underbrace{\sum_{\mathbf{A}\in\{0,1\}^{M\times|\mathcal{N}^{\prime}|}}\!\!\!\!\!\!\!\!\!\!\!\overbrace{\Bigl{(}\xi(\mathcal{N}^{\prime},\mathbf{A})\!\prod_{\upsilon\in\mathcal{N}^{\prime}}\!\!\psi_{\upsilon}(\boldsymbol{a}_{\upsilon})\!\Bigl{)}}^{g_{2}(\mathcal{N}^{\prime},\mathbf{A})}}_{g_{3}(\mathcal{N}^{\prime})}.$ (8)
Here, ${g_{1}(\mathcal{N}^{\prime})}$ denotes the probability of devices in ${\mathcal{N}^{\prime}}$ being active, ${g_{2}(\mathcal{N}^{\prime},\mathbf{A})}$ denotes the probability of successful transmission when devices in ${\mathcal{N}^{\prime}}$ chose ${\mathbf{A}}$, and ${g_{3}(\mathcal{N}^{\prime})}$ denotes the probability of successful transmission for ${\mathcal{N}^{\prime}}$. Lastly, ${\Lambda(\mathbf{\Psi})}$ outputs the probability of successful transmission when considering all possible ${\mathcal{N}^{\prime}}$.
@@
$\underset{\mathbf{\Psi}}{\textrm{maximize}}$ $\Lambda$ (9a)
subject to $\sum_{i=1}^{2^{M}}\psi_{\upsilon}(\boldsymbol{\mathring{a}}_{i})=1\textrm{,}\ \forall\upsilon\in\mathcal{N}^{\prime},$ (9b)
$\mathbf{\Psi}\in[0,1]^{|\mathcal{N}|\times 2^{M}}.$ (9c)
@@ III NEURAL NETWORK-BASED BANDIT
NNBB is a DRL-based RA algorithm that is implemented in each IIoT device. The NNBB components are: (i) an agent (device) that receives a context, (ii) actions (transmission patterns), (iii) an action value ($\psi$) corresponding to each action, and (iv) a reward (feedback signal from the ExC) for updating the action values. The proposed protocol for the alarm transmission is illustrated in Fig. 2. The aforementioned NNBB components, the algorithm implementation, and its computational complexity are explained in detail below.
@@ III-A CONTEXT
An active agent (device) $\upsilon$ observes the context $\boldsymbol{s}_{\upsilon}$. For this, each active agent transmits a pilot signal consisting of $M$ pilot symbols, one on each channel. The BS then receives the signal aggregating pilot signals from all active devices and broadcasts it to agents in the next time slot. The signal $\boldsymbol{s}_{\upsilon}$ received by the active agent $\upsilon$ informs it that some devices are active in the surrounding environment and implicitly captures hidden information about their identity. The goal of the NNBB-based RA mechanism is to make distributed access decisions based on such a context and learning experience.
@@
where $\rho$ represents the average transmit signal-to-noise ratio, ${\boldsymbol{\varphi}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})\in\mathbb{C}^{M\times 1}}$ is the Additive White Gaussian Noise (AWGN) with normalized unit power, and ${\mathbf{c}_{\upsilon}\in\mathbb{C}^{M\times 1}}$ captures the $M$ channel coefficients between the device $\upsilon$ and the BS. After receiving $\boldsymbol{s}$, the BS broadcasts it to devices. Therefore, the context ${\boldsymbol{s}_{\upsilon}\in\mathbb{C}^{M\times 1}}$ received by the active device $\upsilon$ can be written as
@@
where ${\boldsymbol{\hat{\varphi}}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})\in\mathbb{C}^{M\times 1}}$ is the AWGN with normalized power. To simplify the analysis without any loss of generality, we assume that the transmit power of the BS is also $\rho$.
@@ III-B ACTION
After receiving $\boldsymbol{s}_{\upsilon}$, the active agent $\upsilon$ selects a transmission pattern $\boldsymbol{a}_{\upsilon}$, which we shall refer to as its action, and performs a single-hop alarm transmission to the ExC. Recall that with $M$ channels, an active device can choose from $2^{M}$ actions, which are collected as columns of the matrix $\mathbf{\mathring{A}}$. The alarm message consists of a header, an alarm flag, and a payload as shown in Fig. 2. The alarm flag indicates the occurrence of an event, the header carries the metadata, and the payload contains relevant information about the alarm for the ExC.
@@ III-D ACTION VALUE AND SELECTION
An agent aims to maximize its total reward obtained over time. To achieve this, it is attractive to choose those actions that have yielded a reward in the past. This is known as exploitation. However, to discover such reward-accruing actions, an agent must try new actions, which is known as exploration. In our work, an agent balances exploration and exploitation using the $\varepsilon$-greedy method. This method relies on the action value ${q(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i}),\ \forall i=1,\cdots,2^{M}}$, that represents a prediction of the expected reward for agent $\upsilon$ when it takes an action $\boldsymbol{\mathring{a}}_{i}$ with context $\boldsymbol{s}_{\upsilon}$. In the $\varepsilon$-greedy method, an agent $\upsilon$ primarily selects a greedy action (exploitation) ${\boldsymbol{a}_{\upsilon}=\operatorname{argmax}_{\boldsymbol{\mathring{a}}_{i}\in\{0,1\}^{M}}q(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i})}$, with occasional exploration by randomly selecting an action $\boldsymbol{a}_{\upsilon}$ from all possible actions with a probability $\varepsilon$. In this way, every action is eventually selected many times. It is necessary to decrease $\varepsilon$ to gradually emphasize exploitation over exploration. We gradually decrease $\varepsilon$ from $1$ to $0.1$ with a step size of $0.005$ after every alarm event.
@@
Due to channel fading and noise, the number of possible contexts is infinite, leading to an agent encountering a new context every time it becomes active. Thus, maintaining separate action values for each context is not a reasonable approach. Rather, the agent should maintain the action values as a parameterized function and tune its parameters to better match ${q(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i})}$ with the reward obtained after observing the feedback from the ExC. The parameterized function, for which the DNN is employed, allows an estimation of the action value without maintaining a table of action values.
@@
The parameterized action-value function for the agent $\upsilon$ is expressed as ${\hat{q}(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i},\mathbf{w})}$, where $\mathbf{w}$ is a vector of connection weights in DNN layers. The input to the DNN is the received context, as illustrated in Fig. 3. The DNN comprises two fully connected hidden layers and a dense output layer. Notice that two hidden layers are capable of approximating any smooth mapping with arbitrary accuracy [33]. Also, to grapple with the exploding gradients, we use gradient norm clipping [34]. The outputs from the DNN are $2^{M}$ parameterized action values $\hat{q}$. In addition, the DNN employs: (a) ReLU non-linearity, which is a computationally efficient thresholding operation, compared to sigmoid and tanh, which helps alleviate the problem of vanishing gradients, and (b) RMSProp optimizer that maintains a moving average of the squared gradient values for each weight to adjust the update size for each weight. Moreover, the learning rate of the DNN is a hyper-parameter in RMSProp, and we decrease it with a decay rate of $0.015$ per alarm event.
@@
After calculating the reward ${r(\boldsymbol{a}_{\upsilon})}$, the active agent $\upsilon$ stores the tuple ${\{\boldsymbol{s}_{\upsilon},\boldsymbol{a}_{\upsilon},r(\boldsymbol{a}_{\upsilon})\}}$ in its memory buffer $E$. If $E$ is full, it removes an already stored tuple from $E$ in a First-In-First-Out (FIFO) fashion. Indeed, there is no need to store the entire dataset in the agent’s memory. A FIFO memory buffer discards old samples to make room for new ones. This allows an agent to be more reactive to recent changes in its environment.
@@
using RMSProp. Herein, $r_{j}$ and $\hat{q}_{j}$ represent the reward and parameterized action value, respectively, for the $j^{th}$ sample in the mini-batch. Meanwhile, the gradient vector $\nabla_{\mathbf{w}}\Omega$ is clipped as